# Yangqin Jiang, Chao Huang

### v1

@@ 3.1 Experimental Setup
Baseline Methods. Comparisons use two primary model groups: (1) General-purpose vision-capable large models: closed-source ones (GPT series (Hurst et al., 2024): GPT-4o, GPT-5-nano, GPT-5-mini; Gemini family (Comanici et al., 2025): Gemini-1.5-Pro, Gemini-2.5-Pro; Claude family: Claude-3.5-Haiku, Claude-Sonnet-4) and open-source multimodal models (Qwen2.5-VL (Bai et al., 2025), Llama-3.1 (Dubey et al., 2024), GLM series (Hong et al., 2025)); (2) GUI-specialized/fine-tuned models: AutoGLM, AutoGLM-Mobile (Xu et al., 2025), UI-Tars family Qin et al. (2025), V-Droid (Dai et al., 2025), UI-Genie-Agent (Xiao et al., 2025), MobileUse (Li et al., 2025), and two lightly fine-tuned open-source variants (Llama-3.1-8B (ft), GLM-4-9B (ft)).
3.2 Online Agent Evaluation
Using the AndroidLab benchmark, we perform online GUI task evaluations in an Android environment, with results in Table 1. Here, "Ours w/o Cloud LLM" uses only the on-device small model LightAgent, while "Ours w Cloud LLM" refers to the device–cloud collaborative framework where the on-device agent partners with a cloud LLM for task completion. In the "Agent Mode" column, "Simple" means the model directly outputs GUI actions, and "Reason" means it uses a long-horizon reasoning mode. Key findings are as follows:
(i) Small-but-Mighty. The on-device light model LightAgent delivers "small-but-mighty" performance, matching models one size larger (e.g., Qwen2.5-VL-7B-Instruct, GLM-4-9B(ft)) and even some older/lightweight closed-source LLMs (e.g., GPT-5-nano, Claude-3.5-HaiKu, Gemini-1.5-Pro). This stems from our lightweight MLLM training for on-device agents: we first inject GUI-specific knowledge via SFT, then align the training objective with GUI task goals using GRPO. Notably, though LightAgent and GLM-4-9B(ft) share training data, LightAgent —despite being much smaller—does not lag significantly. Additionally, our designed reasoning paradigm helps the small MLLM efficiently use historical context and tap its reasoning capabilities, boosting GUI task performance.
(ii) Favorable Performance-cost Tradeoff. When LightAgent is deployed in a device-cloud setup with a powerful closed-source LLM (e.g., Gemini‑2.5), performance degradation vs. using the closed-source LLM alone is minimal. This approach leverages the on-device model effectively, achieving a strong performance-cost tradeoff. Enabled by our collaborative control framework (combining pre-task complexity assessment and runtime dynamic orchestration), the system adaptively switches between on-device and cloud models based on task difficulty and runtime conditions—maintaining task effectiveness while cutting cloud LLM calls and overall costs.
@@ 3.3 Ablation Study
LightAgent w/o Tuning Reason 3 2.2
LightAgent w/o SFT Reason 12 8.7
LightAgent w/o GRPO Reason 10 7.2
LightAgent w/o Reasoning Simple 12 8.7
LightAgent w/o History Reason 6 4.3
LightAgent Reason 21 15.2
@@
To validate LightAgent ’s introduced techniques, we perform comprehensive ablation studies. Experiments are split into two parts: ablations of on-device model training techniques and of reasoning methods in the overall architecture.
(i) Ablation study for on-device model. As shown in Table 2, we sequentially ablate LightAgent ’s key components. Removing historical context (LightAgent w/o History), GRPO training (LightAgent w/o GRPO), or SFT (LightAgent w/o SFT) each caused significant performance drops—confirming their necessity. Notably, LightAgent w/o SFT outperformed LightAgent w/o GRPO, showing GRPO can independently learn useful policies. This highlights an objective mismatch: while SFT optimizes next-token prediction, GUI tasks demand accurate final actions, which limits optimization for the GUI objective.
We also evaluate GRPO variants (summarized in Figure 5). Two key variants are: LightAgent (small batch), trained with smaller batch sizes (24–32 vs. typical 150–160); and LightAgent-Zero (equivalent to LightAgent w/o SFT), trained from scratch with only GRPO. Figure 5(a) shows larger batches are critical for stable GRPO training—smaller batches cause oscillating task-accuracy rewards and lower success rates. In contrast, LightAgent-Zero’s rewards rise steadily but learn slower, lacking SFT-injected GUI knowledge. Figure 5(b) further illustrates small-batch training leads to highly variable output lengths (especially in reasoning segments), while LightAgent-Zero struggles with stable long-form reasoning—slowing its learning and GUI task performance.
(ii) Ablation study for reasoning methods. As shown in Table 2, LightAgent w/o Reasoning (trained without reasoning segments) shows a marked performance drop vs. the full model. This shows reasoning annotations are critical for smaller models to unlock their potential and gain meaningful capability improvements. Agent mode comparisons in Table 1 offer further insights. Prompting Gemini-2.5-Flash for explicit reasoning significantly improves its GUI performance (SR: 22.5 → 36.2). In contrast, prompting GPT-5-nano for reasoning severely degrades its performance (SR: 18.1 → 2.9). These results reveal a limitation of reasoning techniques: they depend on a model’s baseline capability. For strong models like Gemini-2.5-Flash, reasoning boosts performance; for weaker ones like GPT-5-nano, however, adding reasoning requirements raises task difficulty and impairs results.
3.4 Deep Analysis of Device-Cloud Collaboration
(a) Percentage of steps in device and cloud models.
(b) Cloud steps saved in device-cloud collaboration.
Figure 6: The result of deep analysis of device-cloud collaboration.
@@ 3.4 Deep Analysis of Device-Cloud Collaboration
Figure 6(a) shows the cloud still performs about 65% of steps, reflecting the on-device model’s limited capacity—due to its constrained size—requiring more frequent cloud intervention to sustain task completion rates. Figure 6(b) shows the device‑cloud framework cuts cloud calls by roughly 10%. Notably, more capable cloud models (e.g., GLM-4.5V) exhibit a smaller relative reduction, as they handle a larger share of tasks the on-device model cannot.
3.5 Efficiency Comparison for On-Device Agents
@@
Due to resource constraints on mobile devices, different-sized models show marked differences in on-device runtime efficiency. To quantify this, we assess the response efficiency of three LLMs—our LightAgent (3B), Qwen2.5-VL-7B (7B), and GLM-4.1V-9B (9B)—across two compute setups. Served via vLLM (Kwon et al., 2023), the models are tested on either one NVIDIA RTX 3090 (Single) or two (Double), with all other settings fixed. Results appear in Figure 7; notably, GLM‑4.1V‑9B cannot run on a single RTX 3090 while maintaining required context length, so that configuration’s results are omitted.
@@
On a single RTX 3090, the 7B model’s response time is around 50% longer than our 3B LightAgent ’s, and the 9B model is unusable. With two RTX 3090s, the 7B model remains about 30% slower than the 3B model, and the 9B model becomes usable but has over triple the 3B model’s latency. This confirms model size heavily affects runtime efficiency in resource-constrained environments: our 3B LightAgent, by virtue of its smaller size, delivers a clear efficiency edge while matching larger models in GUI task capabilities. We further note upgrading from one to two RTX 3090s reduces latency by only 15.5% for the 3B model but 27.1% for the 7B model, because the 7B model running near full capacity on a single 3090, amplifying inefficiencies. Thus, under the stricter compute constraints of real-world mobile devices, the efficiency gap between the 7B and 3B models will likely grow.
3.6 Performance on Frequently Used Applications.
@@
To evaluate the agent’s performance on daily mobile apps, we further select common apps, design matching tasks, and report results in Table 3. Experiments show our device-cloud collaborative framework performs well, with no performance degradation compared to a pure cloud-based agent. Moreover, our on-device LightAgent outperforms the larger GLM-4-9B (ft)—even though it underperforms on the original AndroidLab benchmark. We attribute this reversal to our enhanced training pipeline: unlike models fine-tuned directly on raw data, LightAgent is further optimized via GRPO reinforcement learning and augmented with a reasoning paradigm, which boosts the small model’s reasoning ability and generalization. A case of LightAgent ’s performance on TikTok is also reported in Appendix A.6.
4 Related Work
GUI Agent. Autonomous agents show great promise in boosting human task performance. Digital environments have inherently multimodal information (text, images, visual elements), adding complexity and challenges for language models—this has in turn spurred more research on graphical user interface (GUI) agents. Advancements in large-model technologies enable state-of-the-art generalist models (e.g., GPT-5 (OpenAI, 2025), Claude-Sonnet-4 (Anthropic, 2025), Qwen2.5-VL (Bai et al., 2025)) to perform GUI tasks via visual understanding per specific instructions. These models drive progress in visual perception, document parsing, object localization, and reasoning, laying a foundation for multifunctional GUI agents. Meanwhile, many GUI-focused systems (Lu et al., 2025) have emerged from such general models (e.g., UI-Tars (Qin et al., 2025), an end-to-end GUI agent built on Qwen-2-VL (Wang et al., 2024) with strong performance; V-Droid (Dai et al., 2025), which enhances interactive UI element identification by parsing UI state XML and using an agent to verify appropriate actions). Existing GUI agents are typically evaluated for computer and phone use. Phone use introduces extra challenges: stricter on-device compute constraints and operations more reliant on GUI capabilities than MCP (Anthropic, 2024) commands. To tackle these mobile-specific GUI challenges, we propose LightAgent.
Multi-Agent System. As autonomous agent research advances, multi-agent systems are drawing attention, as monolithic approaches struggle with long-context, multimodal scenarios (Belcak et al., 2025). A single agent often fails to handle high-level planning, deep reasoning, and low-level execution. Thus, many studies (Fourney et al., 2024; Wu et al., 2024a; Li et al., 2023) use a coordinator-agent framework: the coordinator interprets user intent and gives instructions, while assistant agents execute tasks, greatly improving complex assignment completion. Systems such as Mobile-Agent-V3 (Ye et al., 2025), CoAct-1 (Song et al., 2025), and Agent-S2 (Agashe et al., 2024) have applied multi-agent architectures to GUI tasks, underscoring collaboration’s importance in addressing complex challenges. Building on multi-agent collaboration advancements, LightAgent introduces a device-cloud collaboration paradigm. It leverages on-device and cloud model strengths to balance cost and performance optimally, better addressing phone-use GUI scenario constraints.
5 Conclusion
In this work, we propose LightAgent —a lightweight device-cloud collaborative framework specifically designed to strike an effective balance between performance and practicality for mobile GUI agents. By engineering a compact yet capable on-device agent and introducing a dynamic orchestration policy, it greatly diminishes reliance on costly cloud models, all while preserving the efficacy of task completion. Comprehensive evaluations demonstrate that LightAgent delivers a favorable trade-off between operational cost and performance, thereby rendering advanced GUI automation more accessible. It marks a meaningful step towards practical and efficient mobile AI agents.
Appendix A Appendix / supplemental material
A.1 LLM Usage Statement
In the preparation of this paper, LLMs are used solely as an auxiliary tool for writing. Their specific role was limited to text polishing—including enhancing the fluency and accuracy of language expression—and assisting in the writing process such as organizing paragraph logic and optimizing sentence structures. LLMs did not participate in research ideation, data analysis, or other core research linkages.
A.2 Instruction Templates
A.2.1 Instruction Template for the On-Device Model.
Template : Instructions for the On-Device Model You are an intelligent agent that performs smartphone tasks by interacting with UI elements labeled with numeric tags.Available Functions 1. tap(index: int) - Tap UI element 2. text(input_str: str) - Insert text (tap field first) 3. long_press(index: int) - Long press UI element 4. swipe(index: int, direction: str, dist: str) - Swipe element - direction: "up", "down", "left", "right" - dist: "short", "medium", "long" 5. back() - Press back button 6. home() - Press home button 7. (interval: int) - Pause (default: 5 seconds) 8. finish(message: str) - Complete taskRequired Output Format <REASONING>
</REASONING>
<STATE_ASSESSMENT>
Current State: [Screen description]
Task Progress: [Completion status]
Next Required Action: [What’s needed]
Expected Outcome: [Action result]
Potential Issues: [Risk considerations]
</STATE_ASSESSMENT>
<CALLED_FUNCTION>
[Single function call only]
</CALLED_FUNCTION>
Guidelines - Execute one action per step - Verify elements exist before interaction - Tap input fields before using text() - Monitor progress to avoid redundant actions - Use finish() only when task complete - Choose direct, efficient paths to completion
A.2.2 Instruction Template for Task Complexity Assessment.
Template : Instructions for Task Complexity Assessment You are an intelligent strategic planning agent that determines the optimal monitoring strategy for smartphone task completion. Your goal is to maximize task completion rate while minimizing cloud model usage costs.Your Role Given a task instruction, determine: 1. When to start monitoring device agent performance 2. How frequently to monitor device agent performance 3. Whether to use cloud model from the beginning for high-risk tasksDevice Agent Capability Assessment Critical Failure Apps (0-15% completion):
- Map.me: Navigation, route planning, complex UI interactions - 0% success rate - PiMusic: Complex music queries, data extraction, multi-screen navigation - 5% success rate High Risk Apps (15-35% completion):
- Financial transactions or data entry
- Navigation and route planning - […] Early Monitoring Required (Start from Step 2-3):
Monitoring Strategy For Critical Failure Apps (0-15% success):
- Start monitoring: Step 1 (immediate)
- Monitoring frequency: Every 2 steps - Consider: Immediate cloud usage for complex tasks For High Risk Apps (15-35% success):
Task-Specific Risk Assessment Analyze the task instruction for these high-risk indicators: 1. Financial/Data Entry: "add transaction", "enter amount", "fill form" 2. Navigation: "find route", "navigate to", "get directions" 3. […]
{Steps Number}
</MONITORING START FROM> <MONITORING FREQUENCY>
{Steps Number}
</MONITORING FREQUENCY>
Guidelines - Prioritize task completion over cost optimization - Use immediate cloud usage for critical failure apps with complex tasks - […]
A.2.3 Instruction Template for Dynamic Orchestration Policy.
Template : Instructions for Dynamic Orchestration Policy You are an intelligent decision-making agent responsible for determining whether to switch from a device model to a cloud model for smartphone task completion. Your primary goal is to maximize task completion success rate.Your Role Analyze the current smartphone screenshot, historical operation information, and task progress to decide if the cloud model should take over from the device model.Decision Criteria Switch to CLOUD model when ANY of the following conditions are met: 1. Immediate Risk Indicators
- Complex Task Pattern: Task involves financial data, navigation, or multi-step forms - Early Failure Signs: Device model shows confusion in first 3 steps - Wrong App Navigation: Device model navigated to completely irrelevant app 2. Progressive Failure Patterns
- Navigation Confusion: Device model appears lost or confused about next steps - Form Completion Issues: Reached correct screen but struggling with form fields - State Misunderstanding: Device model misinterprets current app state or toggle positions 3. Task Progress Assessment
4. Context and Timing Factors
Analysis Framework Immediate Assessment (First 3 Steps):
- Is the device model on the right track?
- Does the current app/screen make sense for the task? - Are there any obvious confusion signs? Progressive Assessment (Steps 4-8):
Critical Decision Points:
- Switch to CLOUD at first sign of struggle
- Prioritize completion over cost - Intervene early rather than late Medium Risk Tasks (Standard Operations):
Low Risk Tasks (Simple Toggles, Basic Navigation):
Output Format After your analysis, output ONLY one of the following decisions: CLOUD - Switch to cloud model (when intervention is needed) DEVICE - Continue with device model (when current approach is working)Guidelines - Prioritize Success: Task completion is more important than cost optimization - Early Intervention: Better to switch too early than too late - Context Awareness: Consider the specific app and task complexity - […] Your analysis should be thorough but your final output must be exactly one word: either "CLOUD" or "DEVICE".
A.2.4 Instruction Template for Reasoning Data Generation.
Template : Instruction Formate for Reasoning Data Generation You are an interface analysis assistant for smartphones. You are provided with a screenshot of a smartphone interface. The interactive elements within the UI are marked with numeric tags starting from 1. For each operable UI element, include the following details: 1. Type of action: Describe the type of interaction available (e.g., navigation, text input, toggle, etc.). 2. Text information: Any visible text associated with the UI element (e.g., labels, placeholders, or descriptions). 3. Action: Summarize what happens when the element is interacted with (e.g., "Tap to navigate to settings," "Toggle to enable/disable Wi-Fi"). 4. State: If the element has a state (e.g., switches for Bluetooth, Wi-Fi), specify whether it is currently "On" or "Off." If no state applies, write "None." 5. Array Indexes: If an element has multiple numeric tags, list all the indexes corresponding to that element. You can call the following functions to interact with those labeled elements to control the smartphone: - [Action Space…]
- Review the current UI elements and their positions - Identify relevant interactive elements for the task 2. History Assessment
- [e…]
3. State Assessment
4. Plan Actions
5. Determine Functions
Your reasoning must explicitly connect your analysis to the function calls you’ll make, ending with the exact function call that matches the provided <CALLED_FUNCTION>.Input Structure You will receive the following input components: 1. Task Instruction
A description of the task to be completed.
2. Screenshot 3. History Info Information about previous states, actions, and their intended goals. 4. Called Function
The specific function you need to justify through your reasoning.
Expected Reasoning Output Example: “‘ <REASONING>
</REASONING>
Your Task Based on the provided task instructions, screenshots, and history information, you will: 1. Analyze the information. 2. Evaluate previous actions and their outcomes. 3. Formulate a clear reasoning process. 4. Ensure your reasoning concludes with and justifies the exact function provided in CALLED FUNCTION. Please output <REASONING>…</REASONING> part with Chain of Thought format step by step.
A.3 Algorithm for GRPO optimization

### v2

@@ 3.2 Online Agent Evaluation
Using the AndroidLab benchmark, we perform online GUI task evaluations in an Android environment, with results in Table 1. Here, "Ours w/o Cloud LLM" uses only the on-device small model OpenPhone, while "Ours w Cloud LLM" refers to the device–cloud collaborative framework where the on-device agent partners with a cloud LLM for task completion. In the "Agent Mode" column, "Simple" means the model directly outputs GUI actions, and "Reason" means it uses a long-horizon reasoning mode. Key findings are as follows:
(i) Small-but-Mighty. The on-device model OpenPhone delivers "small-but-mighty" performance, matching models one size larger (e.g., Qwen2.5-VL-7B-Instruct, GLM-4-9B(ft)) and even some older/lightweight closed-source LLMs (e.g., GPT-5-nano, Claude-3.5-HaiKu, Gemini-1.5-Pro). This stems from our lightweight MLLM training for on-device agents: we first inject GUI-specific knowledge via SFT, then align the training objective with GUI task goals using GRPO. Notably, though OpenPhone and GLM-4-9B(ft) share training data, OpenPhone —despite being much smaller—does not lag significantly. Additionally, our designed reasoning paradigm helps the small MLLM efficiently use historical context and tap its reasoning capabilities, boosting GUI task performance.
(ii) Favorable Performance-cost Tradeoff. When OpenPhone is deployed in a device-cloud setup with a powerful closed-source LLM (e.g., Gemini‑2.5), performance degradation vs. using the closed-source LLM alone is minimal. This approach leverages the on-device model effectively, achieving a strong performance-cost tradeoff. Enabled by our collaborative control framework (combining pre-task complexity assessment and runtime dynamic orchestration), the system adaptively switches between on-device and cloud models based on task difficulty and runtime conditions—maintaining task effectiveness while cutting cloud LLM calls and overall costs.
We also report OpenPhone’s performance on frequently used four applications (i.e., Chrome, TikTok, Reddit, and Gmail) in Appendix A.1.
@@ 3.3 Ablation Study
OpenPhone w/o Tuning Reason 3 2.2
OpenPhone w/o SFT Reason 12 8.7
OpenPhone w/o GRPO Reason 10 7.2
OpenPhone w/o Reasoning Simple 12 8.7
OpenPhone w/o History Reason 6 4.3
OpenPhone Reason 21 15.2
To validate OpenPhone ’s introduced techniques, we perform comprehensive ablation studies. Experiments are split into two parts: ablations of on-device model training techniques and of reasoning methods in the overall architecture. The corresponding results are reported in Table 2, where SN denotes Success Number.
(i) Ablation study for on-device model. As shown in Table 2, we sequentially ablate key components. Removing historical context (OpenPhone w/o History), GRPO training (OpenPhone w/o GRPO), or SFT (OpenPhone w/o SFT) each cause significant performance drops—confirming their necessity. Notably, OpenPhone w/o SFT outperforms OpenPhone w/o GRPO, showing GRPO can independently learn useful policies. This highlights an objective mismatch: while SFT optimizes next-token prediction, GUI tasks demand accurate final actions, limits optimization for the GUI objective.
@@
We also evaluate GRPO variants (Fig. 5): OpenPhone (small batch), trained with reduced batch sizes (24–32 vs. 150–160), and OpenPhone-Zero (equivalent to OpenPhone without SFT), trained from scratch using only GRPO. Fig. 5(a) shows that larger batches stabilize GRPO training, while smaller batches cause oscillating rewards and lower success rates. In contrast, OpenPhone-Zero exhibits steadier reward growth but learns more slowly, lacking prior GUI knowledge from SFT. Additionally, Fig. 5(b) reveals that small-batch training leads to variable output lengths, especially in reasoning sections, whereas OpenPhone-Zero struggles with consistent long-form reasoning, further slowing its learning and GUI performance.
(ii) Ablation study for reasoning methods. As shown in Table 2, OpenPhone w/o Reasoning (trained without reasoning segments) shows a marked performance drop vs. the full model. This shows reasoning annotations are critical for smaller models to unlock their potential and gain meaningful capability improvements. Agent mode comparisons in Table 1 offer further insights. Prompting Gemini-2.5-Flash for explicit reasoning significantly improves its performance (SR: 22.5 → 36.2). In contrast, prompting GPT-5-nano for reasoning severely degrades its performance (SR: 18.1 → 2.9). These results reveal a limitation of reasoning techniques: they depend on a model’s baseline capability. For strong models like Gemini-2.5-Flash, reasoning boosts performance; for weaker ones like GPT-5-nano, however, adding reasoning requirements raises task difficulty and impairs results.
3.4 Deep Analysis of Device-Cloud Collaboration System
(a) $\%$ of steps in device and cloud models.
(b) Cloud steps w device-cloud collaboration.
Figure 6: Analysis result of device-cloud collaboration.
@@ 3.4 Deep Analysis of Device-Cloud Collaboration System
Figure 6(a) shows the cloud still performs about 65% of steps, reflecting the on-device model’s limited capacity—due to its constrained size—requiring more frequent cloud intervention to sustain task completion rates. Figure 6(b) shows the device‑cloud framework cuts cloud calls by roughly 10%. The average number of steps for collaboration here includes the additional model invocation overhead introduced by the collaborative framework. Notably, more capable cloud models (e.g., GLM-4.5V) exhibit a smaller relative reduction, as they handle a larger share of tasks the on-device model cannot.
3.5 Efficiency Comparison
@@
Due to resource constraints on mobile devices, different-sized models show marked differences in on-device runtime efficiency. To quantify this, we assess the response efficiency of three LLMs—our OpenPhone (3B), Qwen2.5-VL-7B (7B), and GLM-4.1V-9B (9B)—across two compute setups. Served via vLLM [14], the models are tested on either one NVIDIA RTX 3090 (Single) or two (Double), with all other settings fixed. Results appear in Figure 7; notably, GLM‑4.1V‑9B cannot run on a single RTX 3090 while maintaining required context length, so that configuration’s results are omitted.
@@
On a single RTX 3090, the 7B model’s response time is approximately 50% longer than that of our 3B OpenPhone, while the 9B model fails to run. Using two RTX 3090s, the 7B model remains about 30% slower than the 3B model, and the 9B model becomes operational but with over triple the latency of the 3B model. This confirms that model size critically impacts runtime efficiency in resource-constrained settings: our compact 3B OpenPhone maintains a clear efficiency advantage while matching the GUI task performance of larger models. We also observe that scaling from one to two GPUs reduces latency by only 15.5% for the 3B model, but by 27.1% for the 7B model, as the latter operates near the capacity limit of a single card, compounding inefficiencies. Consequently, under the stricter computational constraints typical of real-world mobile devices, the efficiency gap between the 7B and 3B models is expected to widen further.
4 Related Work
GUI Agent. Autonomous agents show great promise in boosting human task performance. Digital environments have inherently multimodal information (text, images, visual elements), adding complexity and challenges for language models—this has in turn spurred more research on graphical user interface (GUI) agents. Advancements in large-model technologies enable state-of-the-art generalist models (e.g., GPT-5 [21], Claude-Sonnet-4 [3], Qwen2.5-VL [5], Qwen3-VL [4]) to perform GUI tasks via visual understanding per specific instructions. These models drive progress in visual perception, document parsing, object localization, and reasoning, laying a foundation for multifunctional GUI agents. Meanwhile, many GUI-focused systems [20, 38] have emerged from such general models (e.g., UI-Tars [23], an end-to-end GUI agent built on Qwen-2-VL [28] with strong performance; V-Droid [8], which enhances interactive UI element identification by parsing UI state XML and using an agent to verify appropriate actions). Existing GUI agents are typically evaluated for computer and phone use. Phone use introduces extra challenges: stricter on-device compute constraints and operations more reliant on GUI capabilities than MCP [2] commands. To tackle these mobile-specific GUI challenges, we propose OpenPhone.
Multi-Agent System. As autonomous agent research advances, multi-agent systems are drawing attention, as monolithic approaches struggle with long-context, multimodal scenarios [6]. A single agent often fails to handle high-level planning, deep reasoning, and low-level execution. Thus, many studies [10, 30, 16] use a coordinator-agent framework: the coordinator interprets user intent and gives instructions, while assistant agents execute tasks, greatly improving complex assignment completion. Systems such as Mobile-Agent-V3 [37], CoAct-1 [27], and Agent-S2 [1] have applied multi-agent architectures to GUI tasks, underscoring collaboration’s importance in addressing complex challenges. Building on multi-agent collaboration advancements, OpenPhone introduces a device-cloud collaboration paradigm. It leverages on-device and cloud model strengths to balance cost and performance optimally, better addressing phone-use GUI scenario constraints.
5 Conclusion
In this work, we propose OpenPhone —a lightweight device-cloud collaborative framework specifically designed to strike an effective balance between performance and practicality for mobile GUI agents. By engineering a compact yet capable on-device agent and introducing a dynamic orchestration policy, it greatly diminishes reliance on costly cloud models, all while preserving the efficacy of task completion. Comprehensive evaluations demonstrate that OpenPhone delivers a favorable trade-off between operational cost and performance, thereby rendering advanced GUI automation more accessible. It marks a meaningful step towards practical and efficient mobile AI agents.
Appendix A Appendix / supplemental material
A.1 Performance on Frequently Used Applications.
@@ Appendix A Appendix / supplemental material
To evaluate the agent’s performance on daily mobile apps, we further select common apps, design matching tasks, and report results in Table 3. Experiments show our device-cloud collaborative framework performs well, with no performance degradation compared to a pure cloud-based agent. Moreover, our on-device OpenPhone outperforms the larger GLM-4-9B (ft)—even though it underperforms on the original AndroidLab benchmark. We attribute this reversal to our enhanced training pipeline: unlike models fine-tuned directly on raw data, OpenPhone is further optimized via GRPO reinforcement learning and augmented with a reasoning paradigm, which boosts the small model’s reasoning ability and generalization. We present a case of OpenPhone’s performance on TikTok in Figure 8.
A.2 Experimental Setup
AndroidLab. AndroidLab is an online GUI-task evaluation benchmark built on the Android platform. It comprises nine commonly used applications and 138 evaluation tasks, and supports two input modes: XML mode [33] and SoM (Set-of-Mark) mode [36]. XML mode is tailored to text-only models, where the LLM selects target UI elements directly from the XML representation. SoM mode is intended for multimodal models and uses the Set-of-Mark method: each clickable or focusable element is assigned a unique index, and the LLM specifies elements by their index when issuing operations. Given the growing predominance of MLLMs for GUI tasks, OpenPhone primarily adopts the SoM mode; accordingly, all experimental results reported are obtained using SoM mode.
Additional Frequently Used Applications. Existing academic GUI benchmarks, limited by factors such as reproducibility, often neglect to test many of today’s most widely used mobile applications. To offer a more comprehensive evaluation of OpenPhone, we augmented AndroidLab with four commonly used mobile apps, contributing a total of 25 tasks. The four apps included are Gmail, Chrome, Reddit, and TikTok. Table 4 lists four additional popular mobile apps (Chrome, TikTok, Reddit, and Gmail) and their corresponding tasks; of the 25 tasks in total, only 12 are presented here.
Table 4: Additional App Evaluation Tasks and Descriptions
App Task ID Task Description Evaluation Type
Chrome Browser chrome_2 Set to dark mode operation
TikTok tiktok_1 Go to the homepage of "IShowSpeed" operation
TikTok tiktok_3 Search for videos about "iphone 17" operation
Reddit reddit_1 Join the ChatGPT discussion group operation
Reddit reddit_2 Check the Popular page operation
A.3 Action Space
Table 5: Action Space for Mobile GUI Interaction
Action Parameters Description
Tap index:int Tap the element with the given index number.
Home() none Simulate the home button.
Back() none Simulate the back button.
A.4 Algorithm for GRPO optimization
We report the whole process of GRPO optimization algorithm in Algorithm 2.